Fred is rebuilding his mailbox post. The horizontal portion is 12 inches, where a is 10 inches. The vertical post is 43 inches, where b is 15 inches. Those two pieces meet at a right angle. What is the minimum length necessary for the angled support, c?
1
Expert's answer
2012-01-17T09:37:54-0500
minimum lenght c=sqrt(a*a+b*b)=sqrt(10*10+15*15)=sqrt(100+225)=sqrt(325) ~ 18.03